3. Nonannual compounding period The number of compounding periods in one year is called compounding frequency. The compounding frequency affects both the present and future values of cash flows An investor can invest money with a particular bank and earn a stated interest rate of 13.20%; however, interest will be compounded quarterly. What are the nominal, periodic, and effective interest rates for this investment opportunity? Interest Rates Nominal rate Periodic rate Effective annual rate You want to invest $19,000 and are looking for safe investment options. Your bank is offering you a certificate of deposit that pays a nominal rate of 12% that is compounded quarterly. What is the effective rate of return that you will earn from this investment? o 12.429% 0 12.41096 0 12.55196 0 12.67296
Suppose you decide to deposit $19,000 in a savings account that pays a nominal rate of 11%, but interest is compounded daily. Based on a 365-day year, how much would you have in the account after six months? (Hint: To calculate the number of days, divide the number of months by 12 and multiply by 365.) $20,475.59 $21,278.56 $21,077.82 $20,074.11 О О
Answer #1

a.Nominal rate=13.2%

Periodic rate=(13.2/4)=3.3%

EAR=[(1+APR/m)^m]-1
where m=compounding periods

=(1+0.132/4)^4-1=13.87%(Approx).

b.

EAR=(1+0.12/4)^4-1

=12.551%(Approx).

c.

We use the formula:
A=P(1+r/365)^365n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=19000*(1+0.11/365)^(365*6/12)

=$19000*1.05653186

=$20074.11(Approx).